OpenStudy (anonymous):

[8th Grade Mathematics C] [PLEASE HELP WIll FAN] What is the value of the expression? [(3 + 5) • 2 – 5] • 3

OpenStudy (anonymous):

I also have another that I'm really struggling with.

OpenStudy (anonymous):

but I think the answer would be 18

OpenStudy (haseeb96):

first of all add and subtraction all the terms in brackets and then multiply them u will get the value of that expression

OpenStudy (anonymous):

Ok so would I do it this way 3+5=8 -5=3 3x2=6 then 6x3=13?

OpenStudy (anonymous):

18 I mean

OpenStudy (anonymous):

I believe you but I will wait until they get done replying

OpenStudy (yanasidlinskiy):

@BF4 Heard of PEMDAS Before??? Solve out the parentheses and then the brackets!!:)

OpenStudy (anonymous):

but I will fan you

OpenStudy (yanasidlinskiy):

I'll show you my way:) If you'd like.

OpenStudy (haseeb96):

have u got it how to solve it ?

OpenStudy (mokeira):

BODMAS \[[(3+5) \times 2-5] \times 3\] start with the first bracket (3+5) [8 x 2 - 5] x 3 Now the main bracket. Start with multiplcation [16-5] x 3 Before you multiply make sure you have solved everything in the bracket [16-5] x 3 11 x 3 33 @Haseeb96 am I right?

OpenStudy (anonymous):

Yeah I have heard of PEMDAS and Haseeb I get the last part of how you got it

OpenStudy (anonymous):

Yana show me if you want what ever helps

OpenStudy (haseeb96):

yes u are correct @Mokeira. @BF4 u should write question in this form [ 2.(3+5) -5] .3

OpenStudy (anonymous):

So then the answer is 33?

OpenStudy (yanasidlinskiy):

You gotta follow me though:)!! Here we go:) [(3 + 5) • 2 – 5] • 3 Add the parentheses first:) [8*2-5] *3 Next multiply 8 and 2 [16-5] *3 11*3 = 33

OpenStudy (mokeira):

@BF4 yes it is 33..see my working I have explained how

OpenStudy (yanasidlinskiy):

Does that make sense?

OpenStudy (anonymous):

Ok thanks for helping me @Mokeira and @YanaSidlinskiy

OpenStudy (anonymous):

Oh and almost forgot @Haseeb96

OpenStudy (anonymous):

thanks for helping too!

OpenStudy (yanasidlinskiy):

Welcome!!:)

OpenStudy (yanasidlinskiy):

btw, good luck with ya studies!!:)

OpenStudy (haseeb96):

yes forget me but write yur question in the correct form

OpenStudy (anonymous):

Oh and can you guys answer one more question for me? This one is harder for me. Which placement of grouping symbols in the expression gives it a value of 22? 3 • 4 – 2 + 5 +7 A. 3 • [4 – (2 + 5) + 7] B. (3 • 4) – [2 + (5 + 7)] C. [(3 • 4) – 2] + (5 + 7) D. 3 • (4 – 2) + (5 + 7)

OpenStudy (mokeira):

C..work them all out to confirm

OpenStudy (anonymous):

Thanks for helping all of you